# Limits to scalable evaluation at the frontier: LLM as Judge won’t beat twice the data

### v2

@@ 1.1 Our contributions
In our theoretical work we focus on a standard evaluation setup that encompasses classification, question answering, arena-style comparisons, and safety evaluations. Given an input $x$, a model $m$ produces an output $m(x)$ that receives a binary ground-truth score $s(x,m(x))$. For example, in classification the accuracy score is $s(x,m(x))$ is $1$ if the model’s output $m(x)$ matches the ground-truth label $y(x)$ and 00 otherwise. A judge model provides a binary proxy score $\tilde{s}(x,m(x))$. We imagine that obtaining ground-truth scores is costly and proxy scores are significantly cheaper.
@@
Rather than solely relying on proxy scores, we assume that we have a small number $n$ of samples of ground-truth scores $s$, in addition to a larger number $N$ of proxy scores $\tilde{s}$. A debiasing method is any estimator that provides an unbiased estimate of the true score from the sample of available proxy scores and ground truth scores. In particular, debiasing methods guarantee that ranking models by their estimated performance produces correct model rankings with high probability.
@@ 2 Formal setup and motivation
$b(m)$ =$=\operatorname*{\mathbb{P}}\{s(m)=1\}$
$p(m)$ =$=\operatorname*{\mathbb{P}}\{\tilde{s}(m)=s(m)\mid s(m)=1\}$
$q(m)$ =$=\operatorname*{\mathbb{P}}\{\tilde{s}(m)=s(m)\mid s(m)=0\}$
@@ 2.1 Judge bias can strongly perturb model rankings
We prove Proposition 1 in Appendix C.5. An illustration is provided in Figure 1(a). We use GPT-4 labels to evaluate Claude Sonnet, as well as a set of (ficticious) models on MMLU (Hendrycks et al., 2021). The ficticious models are denoted as ”Claude-Sonnet + $x$” and designed to be strictly better than Claude. This is done by changing wrong predictions of Claude into correct ones until accuracy is improved by $x\%$. Figure 1(b) shows a more realistic setting, with LLama-3 70B judging the top-10 MMLU models according to HELM (Liang et al., 2023). While the ranking is not fully reversed, it is strongly perturbed. This effect can also be observed in prior work: For MT-Bench, the ranking of the best models is not consistently preserved when using LLM judges (Zheng et al., 2024).
@@ 2.2 High agreement rate is insufficient for evaluation
Proposition 2 shows that for agreement $\operatorname*{\mathrm{AG}}(m)=r$, there can be a judge bias of $1-r$ in either direction. This means that even if $\operatorname*{\mathrm{AG}}(m)=r$ was constant across models $m$, we could only reliably rank models for which the true score $\operatorname*{\mathbb{E}}s(m)$ differs by more than $2(1-r)$. As it is common for state-of-the-art models to differ in performance by only low single digit percentages, without further assumptions an agreement rate above $99\%$ (for each evaluated model!) would be required to ensure (asymptotically) correct rankings. For a proof, see Appendix C.6.
@@ 3.1 Uniformly controlling judge bias ensures correct rankings
Let $M$ be a finite set of models such that for any $m,m^{\prime}\in\mathcal{M}$
@@ 3.3 The sample efficiency factor $\tau$
We call $\tau(\hat{\theta})$ the sample efficiency factor of the estimator $\hat{\theta}$. It is our main quantity of interest. In order to ensure it is well-defined, we assume $b(m),p(m),q(m)\in(0,1)$ for the remainder of the text. Our next proposition provides an upper bound on the sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ of PPI, based on the squared Pearson correlation between $s$ and $\tilde{s}$:
@@
3.4 PPI is near-optimal for binary evaluations
The limited sample efficiency gains of PPI compared to the classical estimator raise the question whether we can find an unbiased estimator with lower variance. For black-box estimators that do not model the prompt-conditional distribution $\operatorname*{\mathbb{P}}[(s(m),\tilde{s}(m))|x]$, we show that the answer is no:
@@ 3.4 PPI is near-optimal for binary evaluations
Let $\Theta$ be the set of all unbiased estimators for $\operatorname*{\mathbb{E}}s(m)$ that observe $n$ joint samples $(s(m),\tilde{s}(m))$ and $N$ independent proxy samples $\tilde{s}(m)$. Then, any $\hat{\theta}\in\Theta$ fulfills the variance bound
@@ 3.5 Limited gains for evaluating state-of-the-art models
With Theorem 5 at hand, we can analyze the best sample efficiency factor $\tau$ an unbiased estimator can achieve by making use of model judgments $\tilde{s}$. In this subsection, we focus on the evaluation of frontier models that outperform older models, including the models used to judge them. We use the evaluated model’s score $b(m)$ and the judge’s agreement rate $\operatorname*{\mathrm{AG}}(m)$ as indicators for the respective model’s capabilities. We focus on tasks for which performing and judging the task are of similar difficulty, such that these indicators are comparable. As a prime example, for binary classification, $\operatorname*{\mathrm{AG}}(m)$ equals the judge’s accuracy, while $b(m)$ equals the evaluated model’s accuracy.
@@ 3.6 High agreement rate is insufficient for sample efficiency gains
For any agreement rate $10.5\leq\operatorname*{\mathrm{AG}}(m)<$ 1, there exist values of $b(m),p(m),q(m)\in(0,1)$ such that
@@
We prove the proposition in Appendix C.8. It implies that in some cases despite access to $\tilde{s}$, no unbiased estimator has less variance than $\hat{\theta}^{GT}$. It might be surprising that there are cases in which we cannot make productive use of the proxy evaluations $\tilde{s}$ despite zero judge bias. However, without further assumptions, $\operatorname*{\mathrm{JB}}(m)$ can vary wildly between models, so that we cannot skip estimating it anew for each model. Now, even if $\tilde{s}$ has zero judge bias, the estimator does not ”know” that, but rather has to estimate $\operatorname*{\mathrm{JB}}(m)$ from joint observations $(s,\tilde{s})$. In the configuration from the proposition, estimating the judge bias $\operatorname*{\mathrm{JB}}(m)$ turns out as hard as directly estimating the real score $\operatorname*{\mathbb{E}}s(m)$.
@@ 3.7 Experiments
Our experiments in this section focus on two settings: Multiple-choice question answering on MMLU (Hendrycks et al., 2021) and chatbot evaluation on MT-Bench (Zheng et al., 2024). Additional experiments on TruthfulQA can be found in Appendix B.2. In all cases, we calculate the empirical correlation $\rho(s,\tilde{s})$ between the score $s$ and the proxy $\tilde{s}$ to obtain and plot an upper bound for the sample efficiency factor $\tau_{\max}$ based on Theorem 5.
@@
4 Going beyond binary evaluations.
@@
In terms of analysis, the non-parametric joint likelihood of $(s,\tilde{s})$ resulting from non-binary proxies makes it exceedingly hard to prove an analogon to Theorem 5 for all unbiased estimators. Instead, we focus on the PPI estimator, $\hat{\theta}^{PP}_{\lambda^{*}}$. It is a natural choice of estimator, given its proven optimality in the special case of binary proxies $\tilde{s}$. In order to state a condition similar to $\operatorname*{\mathrm{AG}}\leq b$ from Theorem 6, we assume the proxy $\tilde{s}$ to be bounded in $[0,1]$ and define the soft agreement as
$\operatorname*{\mathrm{SO}}(\tilde{s})\coloneqq\operatorname*{\mathbb{E}}s\tilde{s}+(1-s)(1-\tilde{s}).$
For binary scores $\tilde{s}$, the soft agreement $\operatorname*{\mathrm{SO}}(\tilde{s})$ simply reduces to the agreement $\operatorname*{\mathrm{AG}}(\tilde{s})$. For non-binary proxy scores $\tilde{s}$, we interpret them as probability estimates, such that $\operatorname*{\mathrm{SO}}$ equals the estimated probability of the real score $s$. To make this interpretation valid, we focus on the recalibrated proxy ${R(\tilde{s})=\operatorname*{\mathbb{P}}(s=1|\tilde{s})}$. The following proposition suggests a way to generalize the condition $\operatorname*{\mathrm{AG}}\leq b$:
@@
Proposition 9 states that if the agreement $\operatorname*{\mathrm{AG}}(\tilde{s})$ of a binary proxy $\tilde{s}$ is below the evaluated model’s accuracy $b$, the same is true for the soft accuracy $\operatorname*{\mathrm{SO}}(R(\tilde{s}))$ of the recalibrated proxy $R(\tilde{s})$. Based on this, we use $\operatorname*{\mathrm{SO}}(R(\tilde{s}))\leq b$ to generalize the condition $\operatorname*{\mathrm{AG}}(\tilde{s})\leq b$ from Theorem 6, and prove:
@@ C.2 Theorem 6
We now delve into the details of the proof: Note that the agreement rate is given by $\operatorname*{\mathrm{AG}}(m)=(1-b)q+bq$. We therefore replace $p(m)=\frac{(b-1)q+\operatorname*{\mathrm{AG}}(m)}{b}=\frac{(\delta+x-1)q+x}{b}$ and obtain a formula for $\rho^{2}$ based on $q,x$ and $\delta:$
@@
We next analyze $z_{2}=\frac{\delta x}{\left(2x-1\right)\left(\delta+x-1\right)}$: While $\delta x$ and $12x-$ 1 are positive by assumption, $\delta+x-1$ is negative. Correspondingly, $q=z_{2}$ is not a valid value for $q$ and the maximizer must be one of the other three values.
@@ C.3 Theorem 10
Instead of Theorem 10, we prove a stronger version that allows for the judge to be better than the evaluated model. Concretely, the soft error rate of the judge $(1-\operatorname*{\mathrm{SO}}(R(\tilde{s})))$ is allowed to be smaller than the soft error rate of the evaluated model $1-b$ by a factor $\epsilon<1$. In that case, our upper bound on $\tau$ becomes $\frac{2}{\epsilon}$.
@@
For any proxy score $\tilde{s}$ with $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}}(R(\tilde{s})))$, we have $\rho^{2}(s,\tilde{s})\leq 1-\frac{\epsilon}{2}$. Correspondingly, the sample effficiency of PPI is bounded: $\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq\frac{2}{\epsilon}$.
@@
First, we note that the because the Mean squared error/Brier score is a proper scoring rule (Gneiting & Raftery, 2007), the Bayes-optimal predictor $\operatorname*{\mathrm{R}}(\tilde{s})$ minimizes the mean squared error $\operatorname*{\mathrm{MSE}}(g(\tilde{s}),s)$ for all post-processing functions $g$. At the same time, we have that the squared correlation $\rho^{2}(s,\tilde{s})=R^{2}(s,f(\tilde{s}))=1-\frac{\operatorname*{\mathrm{MSE}}(s,f(\tilde{s}))}{\operatorname*{\mathrm{Var}}(s)}$ falls monotonously in $\operatorname*{\mathrm{MSE}}(s,f(\tilde{s}))$, where $f$ is the affine transformation with the smallest value of $\operatorname*{\mathrm{MSE}}(s,f(\tilde{s}))$ (Larsen & Marx, 2005). This means that the bayes-optimal predictor $\operatorname*{\mathrm{R}}(\tilde{s})$ fulfills $\rho^{2}({s,\operatorname*{\mathrm{R}}(\tilde{s})})\geq\rho^{2}(s,{\tilde{s}})$. As the Bayes-optimal predictor is calibrated, it is sufficient to prove our upper bound for calibrated scores $\tilde{s}$ (i.e. scores $\tilde{s}$ such that $\operatorname*{\mathbb{P}}(s|\tilde{s}=\gamma)=\gamma$ for all $\gamma\in[0,1]$.
@@
where we use that because of calibration, conditional on a prediction $\tilde{s}$, $s$ is one and the squared error is $(1-\tilde{s})^{2}$ with probability $\tilde{s}$. Similarly, the squared error is $\tilde{s}^{2}$ with probability $(1-\tilde{s})$.
@@
[…]
for calibrated scores $\tilde{s}$, as $s=1$ yields a score of $\tilde{s}$ and happens with probability $\tilde{s}$, and $s=0$ happens with probability $(1-\tilde{s})$ and gets a score of $1-\tilde{s}$.
@@
Assume that $b\geq 0.5$. Then for any $0<\epsilon\leq 1$, take any binary proxy $\tilde{s}$ such that $\operatorname*{\mathrm{AG}}\geq 0.5$ and $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{AG}})$. Then, the recalibrated proxy $R(\tilde{s})$ fulfills $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}})$.
We now simply apply Theorem 15. By Proposition 9, the recalibrated version $R(\tilde{s})$ fulfills $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}})$, yielding the desired bound. In the following, we prove Proposition 17:
@@
We first characterize the recalibrated proxy $R(\tilde{s})$ and its relationship with the real score $s$: $R(\tilde{s})$ equals $\operatorname*{\mathbb{P}}(s=1|\tilde{s}=1)$ whenever $\tilde{s}=1$ and $\operatorname*{\mathbb{P}}(s=1|\tilde{s}=0)$ whenever $\tilde{s}=0$. This means that
@@
$\operatorname*{\mathrm{SO}}(R(\tilde{s}))$ $=(bp+(1-b)(1-q))((x_{1})^{2}+(1-x_{1})^{2})$
@@
As $b-1$ is negative, it is sufficient to show that $\frac{\operatorname*{\mathrm{NU}}}{b-1}$ is positive. We take the derivative with respect to $\epsilon$:
@@ C.4 Theorem 12
Assuming $10<b<$ 1 and $\operatorname*{\mathrm{BA}}\neq 0.5$, this is zero if and only if $-(4b\operatorname*{\mathrm{BA}}-2b-2q+1)$ is zero and clearly positive for larger $q$. Correspondingly, as long as the singularities in $\rho^{2}$ lie outside of the $q-$domain, $q=2b\operatorname*{\mathrm{BA}}-b+\frac{1}{2}$ is a unique minimum, while $\rho(s(m),\tilde{s}(m))^{2}$ is maximized either at the minimal or the maximal possible value of $q$.
@@
Next, we assume $\operatorname*{\mathrm{BA}}>\frac{1}{2}$. As both $p$ and $q$ are in $(0,1)$, we get that $q=2\operatorname*{\mathrm{BA}}-p\geq 2\operatorname*{\mathrm{BA}}-1$. Now, the singularities occur at $q=b(2\operatorname*{\mathrm{BA}}-1)$ and $q=2\operatorname*{\mathrm{BA}}b-b+1$, the first of which is smaller than $12\operatorname*{\mathrm{BA}}-$ 1, while the second is larger than $1$. Correspondingly, we can ignore them and insert the minimal $q=2\operatorname*{\mathrm{BA}}-1$. We obtain
@@
These two terms are symmetric in the sense that replacing $b$ in the first term with $1-b$ yields the second term and vice versa. Thus, it is sufficient to maximize the second term with respect to $b$. We obtain
@@
which is negative such that $\rho(s(m),\tilde{s}(m))^{2}$ is maximized at $b=0$, where it equals $12\operatorname*{\mathrm{BA}}-$ 1.
Similarly, for $\operatorname*{\mathrm{BA}}<\frac{1}{2}$, $q$ can range from 00 to $2\operatorname*{\mathrm{BA}}$. Again, the singularities are at $q=b(2\operatorname*{\mathrm{BA}}-1)$ and $q=2\operatorname*{\mathrm{BA}}b-b+1$. This time, the first one is negative, while the second one can easily be seen to be larger than $2\operatorname*{\mathrm{BA}}$. This again allows us to ignore the singularities. Inserting the extreme values yields $\frac{(2\operatorname*{\mathrm{BA}}-1)(b-1)}{2b\operatorname*{\mathrm{BA}}-b+1}$ and $\frac{b(2\operatorname*{\mathrm{BA}}-1)}{2b\operatorname*{\mathrm{BA}}-2\operatorname*{\mathrm{BA}}-b}$, which again are the same after replacing $b$ with $1-b$. Taking the derivative of the first term with respect to $b$ yields

### v3

@@ 1.1 Our contributions
In our theoretical work we focus on a standard evaluation setup that encompasses classification, question answering, arena-style comparisons, and safety evaluations. Given an input $x$, a model $m$ produces an output $m(x)$ that receives a binary ground-truth score $s(x,m(x))$. For example, in classification the accuracy score is $s(x,m(x))$ is $1$ if the model’s output $m(x)$ matches the ground-truth label $y(x)$ and 0 otherwise. A judge model provides a binary proxy score $\tilde{s}(x,m(x))$. We imagine that obtaining ground-truth scores is costly and proxy scores are significantly cheaper.
@@
Rather than solely relying on proxy scores, we assume that we have a small number $n$ of ground-truth scores $s$ sampled IID, in addition to a larger number $N$ of proxy scores $\tilde{s}$. A debiasing method is any estimator that provides an unbiased estimate of the true score from the sample of available proxy scores and ground truth scores. In particular, debiasing methods guarantee that ranking models by their estimated performance produces correct model rankings with high probability.
@@ 2 Formal setup and motivation
$b(m)$ $=\operatorname*{\mathbb{P}}\{s(m)=1\}$
$p(m)$ $=\operatorname*{\mathbb{P}}\{\tilde{s}(m)=s(m)\mid s(m)=1\}$
$q(m)$ $=\operatorname*{\mathbb{P}}\{\tilde{s}(m)=s(m)\mid s(m)=0\}$
@@ 2.1 Judge bias can strongly perturb model rankings
We prove Proposition 1 in Appendix C.5. An illustration is provided in Figure 1(a). We use GPT-4 labels to evaluate Claude Sonnet, as well as a set of (ficticious) models on MMLU (Hendrycks et al., 2021). The ficticious models are denoted as ”Claude-Sonnet + xx” and designed to be strictly better than Claude. This is done by changing wrong predictions of Claude into correct ones until accuracy is improved by $x\%$. Figure 1(b) shows a more realistic setting, with LLama-3 70B judging the top-10 MMLU models according to HELM (Liang et al., 2023). While the ranking is not fully reversed, it is strongly perturbed. This effect can also be observed in prior work: For MT-Bench, the ranking of the best models is not consistently preserved when using LLM judges (Zheng et al., 2024).
@@ 2.2 High agreement rate is insufficient for evaluation
Proposition 2 shows that for agreement $\operatorname*{\mathrm{AG}}(m)=r$, there can be a judge bias of 1r1-r in either direction. This means that even if $\operatorname*{\mathrm{AG}}(m)=r$ was constant across models $m$, we could only reliably rank models for which the true score $\operatorname*{\mathbb{E}}s(m)$ differs by more than 2(1r)2(1-r). As it is common for state-of-the-art models to differ in performance by only low single digit percentages, without further assumptions an agreement rate above $99\%$ (for each evaluated model!) would be required to ensure (asymptotically) correct rankings. For a proof, see Appendix C.6.
@@ 3.1 Uniformly controlling judge bias ensures correct rankings
Let $\mathcal{M}$ be a finite set of models such that for any $m,m^{\prime}\in\mathcal{M}$
@@ 3.3 The sample efficiency factor $\tau$
[…]
@@
3.4 PPI is near-optimal for non-adaptive binary evaluation
The limited sample efficiency gains of PPI compared to the classical estimator raise the question whether we can find an unbiased estimator with lower variance. Given black-box estimators that do not model the prompt-conditional distribution $\operatorname*{\mathbb{P}}[(s(m),\tilde{s}(m))|x]$ and access to $N$ IID samples of proxy scores $\tilde{s}(m)$ and $n$ IID parallel samples $(s(m),\tilde{s}(m))$ we show that the answer is no:
@@ 3.4 PPI is near-optimal for non-adaptive binary evaluation
Let $\Theta$ be the set of all unbiased estimators for $\operatorname*{\mathbb{E}}s(m)$ that observe $n$ joint IID samples $(s(m),\tilde{s}(m))$ and $N$ IID proxy samples $\tilde{s}(m)$ independent of the joint samples. Then, any $\hat{\theta}\in\Theta$ fulfills the variance bound
@@ 3.5 Limited gains for evaluating state-of-the-art models
With Theorem 5 at hand, we can analyze the best sample efficiency factor $\tau$ an unbiased estimator with access to a random sample of $n$ ground truth scores $s(m)$ can achieve by making use of model judgments $\tilde{s}$. In this subsection, we focus on the evaluation of frontier models that outperform older models, including the models used to judge them. We use the evaluated model’s score $b(m)$ and the judge’s agreement rate $\operatorname*{\mathrm{AG}}(m)$ as indicators for the respective model’s capabilities. We focus on tasks for which performing and judging the task are of similar difficulty, such that these indicators are comparable. As a prime example, for binary classification, $\operatorname*{\mathrm{AG}}(m)$ equals the judge’s accuracy, while $b(m)$ equals the evaluated model’s accuracy.
@@ 3.6 High agreement rate is insufficient for sample efficiency gains
For any agreement rate $0.5\leq\operatorname*{\mathrm{AG}}(m)<1$, there exist values of $b(m),p(m),q(m)\in(0,1)$ such that
@@
We prove the proposition in Appendix C.8. It implies that in some cases despite access to $\tilde{s}$, no unbiased estimator has less variance than $\hat{\theta}^{GT}$. It might be surprising that there are cases in which we cannot make productive use of the proxy evaluations $\tilde{s}$ despite zero judge bias. However, without further assumptions, $\operatorname*{\mathrm{JB}}(m)$ can vary wildly between models, so that we cannot skip estimating it anew for each model. Now, even if $\tilde{s}$ has zero judge bias, the estimator does not ”know” that, but rather has to estimate $\operatorname*{\mathrm{JB}}(m)$ from joint observations (s,s~)($\tilde{s}$). In the configuration from the proposition, estimating the judge bias $\operatorname*{\mathrm{JB}}(m)$ turns out as hard as directly estimating the real score $\operatorname*{\mathbb{E}}s(m)$.
@@ 3.7 Experiments
[…]
@@
4 Going beyond binary evaluation
@@
In terms of analysis, the non-parametric joint likelihood of (s,s~)($\tilde{s}$) resulting from non-binary proxies makes it exceedingly hard to prove an analogon to Theorem 5 for all unbiased estimators. Instead, we focus on the PPI estimator, $\hat{\theta}^{PP}_{\lambda^{*}}$. It is a natural choice of estimator, given its proven optimality in the special case of binary proxies $\tilde{s}$. In order to state a condition similar to $\operatorname*{\mathrm{AG}}\leq b$ from Theorem 6, we assume the proxy $\tilde{s}$ to be bounded in $[0,1]$ and define the soft agreement as
$(1s~).\operatorname*{\mathrm{SO}}(\tilde{s})\coloneqq\operatorname*{\mathbb{E}}s\tilde{s}$$(1-\tilde{s})$.
[…]
@@
[…]
@@ C.2 Theorem 6
We now delve into the details of the proof: Note that the agreement rate is given by $\operatorname*{\mathrm{AG}}(m)=(1-b)q+bp$. We therefore replace $p(m)=\frac{(b-1)q+\operatorname*{\mathrm{AG}}(m)}{b}=\frac{(\delta+x-1)q+x}{b}$ and obtain a formula for $\rho^{2}$ based on $q,x$ and $\delta:$
@@
We next analyze $z_{2}=\frac{\delta x}{\left(2x-1\right)\left(\delta+x-1\right)}$: While $\delta x$ and 2x12x-1 are positive by assumption, $\delta+x-1$ is negative. Correspondingly, $q=z_{2}$ is not a valid value for $q$ and the maximizer must be one of the other three values.
@@ C.3 Theorem 10
[…]
@@
[…]
@@
[…]
@@
where we use that because of calibration, conditional on a prediction $\tilde{s}$, $s$ is one and the squared error is $(1-\tilde{s})^{2}$ with probability $\tilde{s}$. Similarly, the squared error is $\tilde{s}^{2}$ with probability (1s~)(1-$\tilde${s}).
@@
$\operatorname*{\mathrm{SO}}=\int_{0}^{1}(\tilde{s}^{2}+(1-\tilde{s})^{2})p(\tilde{s})d\tilde{s}=\int_{0}^{1}(2\tilde{s}(\tilde{s}-1)+1)p(\tilde{s})d\tilde{s}$
for calibrated scores $\tilde{s}$, as $s=1$ yields a score of $\tilde{s}$ and happens with probability $\tilde{s}$, and $s=0$ happens with probability (1s~)(1-$\tilde${s}) and gets a score of $1-\tilde{s}$.
@@
Assume that $b\geq 0.5$. Then for any $0<\epsilon\leq 1$, take any binary proxy $\tilde{s}$ such that $\operatorname*{\mathrm{AG}}\geq 0.5$ and $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{AG}})$. Then, the recalibrated proxy R(s~)R($\tilde${s}) fulfills $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}})$.
We now simply apply Theorem 15. By Proposition 9, the recalibrated version R(s~)R($\tilde${s}) fulfills $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}})$, yielding the desired bound. In the following, we prove Proposition 17:
@@
We first characterize the recalibrated proxy R(s~)R($\tilde${s}) and its relationship with the real score $s$: R(s~)R($\tilde${s}) equals $\operatorname*{\mathbb{P}}(s=1|\tilde{s}=1)$ whenever $\tilde{s}=1$ and $\operatorname*{\mathbb{P}}(s=1|\tilde{s}=0)$ whenever $\tilde{s}=0$. This means that
@@
$)\displaystyle\operatorname*{\mathrm{SO}}(R(\tilde{s}$)) $=(bp+(1-b)(1-q))((x_{1})^{2}+(1-x_{1})^{2})$
@@
As b1b-1 is negative, it is sufficient to show that $\frac{\operatorname*{\mathrm{NU}}}{b-1}$ is positive. We take the derivative with respect to $\epsilon$:
@@ C.4 Theorem 12
Assuming $0<b<1$ and $\operatorname*{\mathrm{BA}}\neq 0.5$, this is zero if and only if $(4b\operatorname*{\mathrm{BA}}-2b-2q+1)$ is zero and clearly positive for larger $q$. Correspondingly, as long as the singularities in $\rho^{2}$ lie outside of the $q$-domain, $q=2b\operatorname*{\mathrm{BA}}-b+\frac{1}{2}$ is a unique minimum, while $\rho(s(m),\tilde{s}(m))^{2}$ is maximized either at the minimal or the maximal possible value of $q$.
@@
Next, we assume $\operatorname*{\mathrm{BA}}>\frac{1}{2}$. As both $p$ and $q$ are in $(0,1)$, we get that $q=2\operatorname*{\mathrm{BA}}-p\geq 2\operatorname*{\mathrm{BA}}-1$. Now, the singularities occur at $q=b(2\operatorname*{\mathrm{BA}}-1)$ and $q=2\operatorname*{\mathrm{BA}}b-b+1$, the first of which is smaller than $2\operatorname*{\mathrm{BA}}-1$, while the second is larger than $1$. Correspondingly, we can ignore them and insert the minimal $q=2\operatorname*{\mathrm{BA}}-1$. We obtain
@@
These two terms are symmetric in the sense that replacing $b$ in the first term with 1b1-b yields the second term and vice versa. Thus, it is sufficient to maximize the second term with respect to $b$. We obtain
@@
which is negative such that $\rho(s(m),\tilde{s}(m))^{2}$ is maximized at $b=0$, where it equals $2\operatorname*{\mathrm{BA}}-1$.
Similarly, for $\operatorname*{\mathrm{BA}}<\frac{1}{2}$, $q$ can range from 0 to $2\operatorname*{\mathrm{BA}}$. Again, the singularities are at $q=b(2\operatorname*{\mathrm{BA}}-1)$ and $q=2\operatorname*{\mathrm{BA}}b-b+1$. This time, the first one is negative, while the second one can easily be seen to be larger than $2\operatorname*{\mathrm{BA}}$. This again allows us to ignore the singularities. Inserting the extreme values yields $\frac{(2\operatorname*{\mathrm{BA}}-1)(b-1)}{2b\operatorname*{\mathrm{BA}}-b+1}$ and $\frac{b(2\operatorname*{\mathrm{BA}}-1)}{2b\operatorname*{\mathrm{BA}}-2\operatorname*{\mathrm{BA}}-b}$, which again are the same after replacing $b$ with 1b1-b. Taking the derivative of the first term with respect to $b$ yields